OpenStudy (anonymous):

Somebody please help!!! Solve this problem and put it into slope-intercept form. x - 4y = 2 which is y = 1/4x - 1/2 Now I know that this basically means that the slope is 1/4 or 1 over 4 but how do I graph the -1/2 part?

Directrix (directrix):

Are you working this problem: y = 1/4x - 1/2 or the one above it?

OpenStudy (anonymous):

That is the problem above but put into slope-intercept form. So yeah, I'm working with y= 1/4 x+ −1/2

OpenStudy (anonymous):

y = 1/4x - 1/2

Directrix (directrix):

This --> y = 1/4x - 1/2 is not what you get when you solve for y on this -->x + 2y = 8

Directrix (directrix):

x + 2y = 8 2y + x = 8 2y = -x + 8 y = (-1/2) * x + 4 is what I get when solving for y in the top equation.

OpenStudy (anonymous):

Whoops. :/ Mistake on my part. The equation is supposed to be x - 4y = 2